Mass vs T 1.6 1.4 y = 2.2486x0.41 1.2 1 Period (s) 0.8 0.6 0.4 0.2 0 0 0.05 0.1 0.15 0.2 0.25 0.3 0.35 Mass (kg)

how this mathematical model relates to the physical hypothesis given in Eq. 2.

т T= 21 kequation 2

Calculate the spring stiffness, k, from your fit parameter, A.

“y = A*sqrt(x).”
Answer #1

Answer :-- the time period can be written as

T = \frac{2\pi}{\sqrt{k}}\sqrt{m}...........................................(1)

comparing with given equation as ploted by the experiment is we have

  • \frac{2\pi}{\sqrt{k}} = 2.2486
  • k= \frac{4\times 3.14 \times 3.14}{2.2486 \times 2.2486}
  • k=7.8000049 ...........................(answer)